# Thermoelectric effects of an interacting hadron gas in the presence of an external magnetic field

### v1

@@ II.1.4 Repulsive Mean-Field Hadron Resonance Gas Model (RMFHRG)
Here, to simplify the further calculation, we can choose the magnetic field along the $z$ direction. The direction of the electric field and the temperature gradient are perpendicular to the $z$ axis i.e. it is in the $x-y$ plane. Under these conditions, the components of the electric current in the $x-y$ plane are given as,
@@
$(-)\frac{\partial f_{i}^{0}}{\partial\omega_{i}}-\sum_{i}\frac{g_{i}q_{i}}{3T}\int\frac{d^{3}k_{i}}{(2\pi)^{3}}\frac{v_{i}^{2}\tau^{i}_{R}\left(\omega_{i}-b_{i}h\right)}{1+(\Omega_{c_{i}}\tau)^{2}}\bigg{]}$
@@
Eq.(76) and (77) can be written in a compact form by introducing the following integrals,
@@
The integrals as given in Eq.(78)-Eq.(81) allows us to write Eq.(76) and Eq.(77), respectively, as
@@
Here, in the presence of a magnetic field, the magneto-Seebeck coefficient ($S_{B}$) can be determined by setting $j_{x}=j_{y}=0$ so that the electric field becomes proportional to the temperature gradient. For $j_{x}=0$ and $j_{y}=0$ we can solve Eq.(82) and (83) to get $E_{x}$ and $E_{y}$ in terms of temperature gradients $\frac{dT}{dx}$ and $\frac{dT}{dy}$ as,
@@
Eq.(84) and (85) can be written in a compact form in the following way,
@@
In the above calculation, for the sake of simplicity, we take the thermal averaged relaxation time after integrating energy-dependent relaxation time over the equilibrium distribution function. The thermal averaged relaxation time ($\tau_{R}^{i}$) for ith hadron species can be expressed in terms of scattering cross-section as Das:2020beh ,
@@
here, $\sigma=4\pi r_{h}^{2}$ is the total scattering cross-section for the hard spheres, and it is independent of both temperature and baryon chemical potential. $\mathcal{K}_{1},\mathcal{K}_{2}$ are modified Bessel functions of the first and second order.
@@ III Results and Discussions
We take the different versions of hadron resonance gas models with discrete particle spectrum, including all the hadrons and their resonances up to the mass cutoff $\Lambda$ = 2.6 GeV ParticleDataGroup:2008zun . The hard-scattering approximation is used to estimate the relaxation time of the hadrons. Here, we consider a uniform radius, $r_{m}$ = 0.2 fm for mesons and $r_{b,\bar{b}}$ = 0.62 fm for baryons and anti-baryons Pradhan:2022gbm . For the VDWHRG model, the van der Waals parameters are taken as $a$ = 0.926 GeV fm3 and $b$ = (16/3)$\pi r^{3}$ Pradhan:2022gbm ; Samanta:2017yhh . In this section, we present the results for different thermoelectric coefficients of a hot and dense hadron gas both in the absence and presence of a nonzero external magnetic field.
@@ III.1 In the absence of magnetic field
Figure 2: Comparison of Seebeck coefficient ($S$) in different models as a function of temperature at a baryonic chemical potential $\mu_{B}$ = 0.10 GeV. The dotted magenta line shows the results obtained in ideal HRG in Ref. Das:2020beh , and the dashed blue line is obtained in Nambu–Jona Lasinio (NJL) model Abhisek2022 .
@@
Fig. 1 shows the Seebeck coefficient, $S$ as a function of temperature for three different baryon chemical potentials at 0.10, 0.30, and 0.40 GeV, in the absence of an external magnetic field. The Seebeck coefficient obtained in the IHRG model is shown by a black solid line, whereas the red dotted line represents the EVHRG model. The green and cyan dashed lines represent the results for the RMFHRG and VDWHRG models, respectively. Here, we observe that for the given range of temperature, $S$ is negative for a hot hadron gas, and is universally decreasing with an increase in temperature. However, the degree of decrease depends on the model. The heat flow for a relativistic system can only be defined for any conserved current. For the case of hadron gas, the heat current corresponds to the net baryon current. It is to be noted that baryons have a dominant contribution to the Seebeck coefficient, whereas the contribution of mesons comes only through the enthalpy of the system. For the case of HRG models, the entropy production for lighter mesons like pions, kaons, etc., is substantial. Therefore, the enthalpy per baryon ($h$) exceeds the single particle energy ($\omega_{i}$), hence the term ($\omega_{i}-b_{i}h$) in $\mathcal{I}_{1}/T^{2}$ becomes negative in Eq. (51), and as a result, $S$ becomes negative. The negative sign of $S$ indicates that the induced electric field in the medium is aligned opposite to the temperature gradient. For the condensed matter systems, $S$ can be both positive or negative corresponding to the contribution of holes and electrons, respectively. In our case, we observe that the scaled electrical conductivity, $\sigma_{el}/T$, decreases as a function of temperature, in line with what was observed in earlier studies Das:2020beh ; Pradhan:2022gbm ; Singh:2023pwf ; Das:2019wjg . Therefore, the magnitude of the Seebeck coefficient increases as a function of temperature but in a negative direction. As shown in the left panel of Fig. 1, results from all the models coincide up to a temperature $T\approx 0.14$ GeV, after which deviations are observed. The effect of interactions between the hadrons seems to be prominent after this temperature. One can observe from Fig. 1 that as $\mu_{B}$ increases, $S$ increases for all the models and the deviation from each model starts appearing even at low temperatures; around $T\approx 0.12$ GeV for $\mu_{B}=$ 0.30 GeV and $T\approx 0.11$ GeV for $\mu_{B}=$ 0.40 GeV.
@@
In Fig. 2, we compare our results of the Seebeck coefficient with those obtained in different phenomenological works at $\mu_{B}$ = 0.10 GeV. The $S$ obtained in Ref. Das:2020beh using the IHRG model (Magenta dotted line) is found to be less in magnitude as compared to our results. This can be attributed to the use of different hadronic radii for the hard-sphere scattering cross-section to estimate the relaxation time. In Ref. Das:2020beh , a uniform radius, $r_{h}$ = 0.3 fm, is taken for all the hadrons, whereas we choose $r_{M}$ = 0.2 fm for mesons and $r_{B}$ = 0.62 fm for (anti)baryons, which give a better description for mesons and baryons. The result from the Nambu–Jona Lasinio (NJL) model Abhisek2022 for the confined phase is shown in a dotted dashed blue line.
@@
Fig. 3 represents the Thomson coefficient, $Th$ as a function of temperature for the same baryon chemical potentials, i.e., at $\mu_{B}$ = 0.10, 0.30, and 0.40 GeV. In condensed matter systems, the Thomson coefficient for a system refers to the absorption or release of heat ($Q$) in the system when a current $\vec{j}$ flows through it corresponding to the temperature gradients $\vec{\nabla}T$, as $Q=Th~{}\vec{j}\cdot\vec{\nabla}T$ ioffe1957physics . The positive (negative) values of the Thomson coefficient lead to the absorption (release) of the heat in the system. Therefore, the presence of a non-zero Thomson coefficient may affect the cooling rate of the system. From fig. 3 we observe that similar to the Seebeck coefficient, Th also gives negative values for all three cases, decreasing universally as a function of temperature. As given in Eq. (54), the Thomson coefficient is also defined as the product of temperature to the rate of change of the Seebeck coefficient with respect to temperature. So, higher is the steepness of the Seebeck coefficient in Fig. 1, the higher is the value of $Th$ in magnitude as a function of $T$. Therefore, in a similar behaviour to the Seebeck coefficient, the value of $Th$ is found to be more negative (i.e., more amount of heat is released) towards high temperature for EVHRG and VDWHRG. Whereas all these models overlap at lower values of temperature nearly up to $\approx 0.14$ GeV for the case of $\mu_{B}$ = 0.10 GeV. As the value of $\mu_{B}$ increases, the models start deviating from each other even at low temperatures, as in the case of the Seebeck coefficient, which is explained above.
@@
To study the efficiency of QGP medium as a thermoelectric medium, we draw the thermoelectric figure of merit, $ZT$ as a function of temperature in Fig. 4. As given in Eq. (59), $ZT$ can be defined as the ratio of $S^{2}$ to the Lorenz number ($L$), which is proportional to the ratio of thermal to electrical conductivity and defined as $L=\kappa_{0}/\sigma_{el}T$ Pradhan:2022gbm ; Singh:2023pwf . Since it is proportional to $S^{2}$, we observed that $ZT$ increases both as a function of $T$ and $\mu_{B}$. The rise in $\mu_{B}$ gives rise to $ZT$ because of the presence of more baryons as compared to antibaryons. Hence, the net thermoelectric current increases, which ultimately improves the thermoelectric efficiency of the medium. The sudden increase of ZT for the VDWHRG case can be attributed to the rise of pressure as well as number density, which subsequently affects other transport properties, due to the interplay between both attractive and repulsive interactions between the hadrons. However, the decreasing trend in the VDWHRG case may be due to the effect of liquid gas phase transition, which comes at a lower temperature with increasing baryochemical potential.
@@ III.2 In presence of magnetic field
Here, we discuss the magneto-Seebeck coefficient, $S_{B}$ in the presence of a non-zero magnetic field as a function of $T$ for different values of $\mu_{B}$. The upper panels of Fig. 5 represent $S_{B}$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at magnetic field 0.1 $m_{\pi}^{2}$, whereas the lower panels represent those for the magnetic field 1.0 $m_{\pi}^{2}$. For the case of $eB$ = 0.1 $m_{\pi}^{2}$, we observe that $S_{B}$ increases with the temperature almost linearly up to 0.1 GeV and then starts to decrease at higher temperatures. Similar to what was observed in the case of a vanishing magnetic field, all the models agree up to a certain temperature and then start deviating from each other at higher temperatures. Furthermore, we also observe that the values of $S_{B}$ increase with the increasing values of $\mu_{B}$. In contrast to the zero magnetic field case, as shown in Eq. (87), $S_{B}$ has a dependence on the Hall-like components of electrical conductivity and other integral terms which comes into picture because of the presence of the external magnetic field. The Hall-like component of electrical conductivity, $\sigma_{H}/T$, increases with $\mu_{B}$ because of an increase in net baryonic contributions. It is also to be noted here that there is no net contribution of mesons in the integral $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$, but it enters through enthalpy per baryon ($h$). The effect of the magnetic field is significant on $\mathcal{I}_{31}/T^{2}$ at lower temperatures, but for higher temperatures, it is not effective due to the lower value of relaxation time. We also observe that $\mathcal{I}_{42}/T^{2}$ first increases with temperature and then starts to decrease. Higher is the magnetic field, lower is the $\mathcal{I}_{42}/T^{2}$. The order of magnitude of $\sigma_{el}/T$ is larger than $\sigma_{H}/T$, whereas the order of magnitude of $\mathcal{I}_{31}/T^{2}$ is close to $\mathcal{I}_{42}/T^{2}$. Hence, the behaviour of the magneto-Seebeck coefficient can be approximated as the ratio of $(\mathcal{I}_{31}/T^{2})/(\sigma_{el}/T)$ which can be seen in Fig. 5. The dependency of $\sigma_{el}/T$ on $T$ and $\mu_{B}$ in the presence of a magnetic field results in first increase of $S_{B}$ at low $T$ and then decrease at high $T$. This is because, in the presence of a magnetic field, the higher value of $\Omega_{c_{i}}\tau_{R_{i}}$ at the lower temperature range affects the mesonic contribution to $\sigma_{el}/T$ significantly. This decrease in mesonic contribution gets compensated for by an increase in baryon contribution with increasing $\mu_{B}$. On the other hand, this decrease in the mesonic contribution is not compensated by those increasing baryonic contributions for higher temperature ranges. Hence, $\sigma_{el}/T$ shows decreasing trend with increasing $\mu_{B}$ for higher temperatures. For a higher value of the magnetic field as shown in the lower panels in Fig. 5, we observe a minimum for $S_{B}$ while going to the high $\mu_{B}$ region. This minimum shifts towards high $T$ when $\mu_{B}$ increases further.
@@
Finally, we discuss the variation of the figure of merit with temperature and baryon chemical potential in the presence of an external magnetic field. Fig. 7 shows $ZT$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at two different magnetic field orders, $eB$ = 0.1 $m_{\pi}^{2}$ in the upper panels, whereas the lower panels represent those for $eB$ = 1.0 $m_{\pi}^{2}$. In the presence of a magnetic field, the behaviour of $ZT$ is different at low temperatures as compared to those in the case of vanishing magnetic fields. At higher temperatures, the results are nearly the same as in the case of the vanishing magnetic field, as shown in Fig. 4. As the value of $\mu_{B}$ increases, a minimum appears at low temperatures. For the case of $eB$ = 1.0 $m_{\pi}^{2}$, $ZT$ gets fully affected throughout the temperature range considered. This is because $ZT$ is proportional to the square of the magneto-Seebeck coefficient, and hence, the effect of the magnetic field on $S_{B}$ is reflected in $ZT$. The presence of a magnetic field results in the highly non-monotonic behaviour of $ZT$ as a function of both $T$ and $\mu_{B}$. We observe that the values of the thermoelectric figure of merit overlap for all the models at low temperatures, but a large deviation is observed at high temperatures. Along with the Ohmic-like component of thermal conductivity ($\kappa_{0}$), the Hall-like component of thermal conductivity ($\kappa_{H}$ingh:2023pwf also contributes to $ZT$ in the presence of an external magnetic field. However, the effect of $\kappa_{H}$ becomes more significant in the presence of strong magnetic field.

### v2

@@ II.1.4 Repulsive mean-field hadron resonance gas model
Here, to simplify the calculation, we can choose the magnetic field along the $z$ direction. The direction of the electric field and the temperature gradient are perpendicular to the $z$ axis i.e. it is in the $x-y$ plane. Under these conditions, the components of the electric current in the $x-y$ plane are given as,
@@
Equations (76) and (77) can be written in a compact form by introducing the following integrals,
@@
The integrals as given in Eqs.(78)-(81) allows us to write Eq.(76) and Eq.(77), respectively, as
@@
Here, in the presence of a magnetic field, the magneto-Seebeck coefficient ($S_{B}$) can be determined by setting $j_{x}=j_{y}=0$ so that the electric field becomes proportional to the temperature gradient. For $j_{x}=0$ and $j_{y}=0$ we can solve Eqs.(82) and (83) to get $E_{x}$ and $E_{y}$ in terms of temperature gradients $\frac{dT}{dx}$ and $\frac{dT}{dy}$ as,
@@
Equations (84) and (85) can be written in a compact form in the following way,
@@
In the above calculation, for the sake of simplicity, we take the thermal averaged relaxation time after integrating energy-dependent relaxation time over the equilibrium distribution function. The thermal averaged relaxation time ($\tau_{R}^{i}$) for $i$th hadron species can be expressed in terms of scattering cross section as Das:2020beh ,
@@
here, $\sigma=4\pi r_{h}^{2}$ is the total scattering cross section for the hard spheres, and it is independent of both temperature and baryon chemical potential. $\mathcal{K}_{1},\mathcal{K}_{2}$ are modified Bessel functions of the first and second order.
@@ III Results and Discussions
We take the different versions of hadron resonance gas models with discrete particle spectrum, including all the hadrons and their resonances up to the mass cutoff $\Lambda$ = 2.6 GeV ParticleDataGroup:2008zun . The hard-scattering approximation is used to estimate the relaxation time of the hadrons. Here, we consider a uniform radius, $r_{m}$ = 0.2 fm for mesons and $r_{b,\bar{b}}$ = 0.62 fm for baryons and antibaryons Pradhan:2022gbm . For the VDWHRG model, the van der Waals parameters are taken as $a$ = 0.926 GeV fm3 and $b$ = (16/3)$\pi r^{3}$ Pradhan:2022gbm ; Samanta:2017yhh . In this section, we present the results for different thermoelectric coefficients of a hot and dense hadron gas both in the absence and presence of a nonzero external magnetic field.
@@ III.1 In the absence of magnetic field
Figure 2: Comparison of Seebeck coefficient ($S$) in different models as a function of temperature at a baryonic chemical potential $\mu_{B}$ = 0.10 GeV. The dotted magenta line shows the results obtained in ideal HRG in Ref. Das:2020beh , and the dashed blue line is obtained in the Nambu–Jona Lasinio (NJL) model Abhisek2022 .
@@
Figure 1 shows the Seebeck coefficient, $S$ as a function of temperature for three different baryon chemical potentials at 0.10, 0.30, and 0.40 GeV, in the absence of an external magnetic field. The Seebeck coefficient obtained in the IHRG model is shown by a black solid line, whereas the red dotted line represents the EVHRG model. The green and cyan dashed lines represent the results for the RMFHRG and VDWHRG models, respectively. Here, we observe that for the given range of temperature, $S$ is negative for a hot hadron gas, and is universally decreasing with an increase in temperature. However, the degree of decrease depends on the model. The heat flow for a relativistic system can only be defined for any conserved current. For the case of hadron gas, the heat current corresponds to the net baryon current. It is to be noted that baryons have a dominant contribution to the Seebeck coefficient, whereas the contribution of mesons comes only through the enthalpy of the system. For the case of HRG models, the entropy production for lighter mesons, such as pions and kaons, is substantial. Therefore, the enthalpy per baryon ($h$) exceeds the single particle energy ($\omega_{i}$), hence the term ($\omega_{i}-b_{i}h$) in $\mathcal{I}_{1}/T^{2}$ becomes negative in Eq. (51), and as a result, $S$ becomes negative. The negative sign of $S$ indicates that the induced electric field in the medium is aligned opposite to the temperature gradient. For the condensed matter systems, $S$ can be both positive or negative corresponding to the contribution of holes and electrons, respectively. In our case, we observe that the scaled electrical conductivity, $\sigma_{el}/T$, decreases as a function of temperature, in line with what was observed in earlier studies Das:2020beh ; Pradhan:2022gbm ; Singh:2023pwf ; Das:2019wjg . Therefore, the magnitude of the Seebeck coefficient increases as a function of temperature but in a negative direction. As shown in the left panel of Fig. 1, results from all the models coincide up to a temperature $T\approx 0.14$ GeV, after which deviations are observed. The effect of interactions between the hadrons seems to be prominent after this temperature. One can observe from Fig. 1 that as $\mu_{B}$ increases, $S$ increases for all the models and the deviation from each model starts appearing even at low temperatures; around $T\approx 0.12$ GeV for $\mu_{B}=$ 0.30 GeV and $T\approx 0.11$ GeV for $\mu_{B}=$ 0.40 GeV.
@@
In Fig. 2, we compare our results of the Seebeck coefficient with those obtained in different phenomenological works at $\mu_{B}$ = 0.10 GeV. The $S$ obtained in Ref. Das:2020beh using the IHRG model (Magenta dotted line) is found to be less in magnitude as compared to our results. This can be attributed to the use of different hadronic radii for the hard-sphere scattering cross section to estimate the relaxation time. In Ref. Das:2020beh , a uniform radius, $r_{h}$ = 0.3 fm, is taken for all the hadrons, whereas we choose $r_{M}$ = 0.2 fm for mesons and $r_{B}$ = 0.62 fm for (anti)baryons, which give a better description for mesons and baryons. The result from the Nambu–Jona Lasinio (NJL) model Abhisek2022 for the confined phase is shown in a dotted dashed blue line.
@@
Fig. 3 represents the Thomson coefficient, $Th$ as a function of temperature for the same baryon chemical potentials, i.e., at $\mu_{B}$ = 0.10, 0.30, and 0.40 GeV. In condensed matter systems, the Thomson coefficient for a system refers to the absorption or release of heat ($Q$) in the system when a current $\vec{j}$ flows through it corresponding to the temperature gradients $\vec{\nabla}T$, as $Q=Th~{}\vec{j}\cdot\vec{\nabla}T$ ioffe1957physics . The positive (negative) values of the Thomson coefficient lead to the absorption (release) of the heat in the system. Therefore, the presence of a nonzero Thomson coefficient may affect the cooling rate of the system. From fig. 3 we observe that similar to the Seebeck coefficient, Th also gives negative values for all three cases, decreasing universally as a function of temperature. As given in Eq. (54), the Thomson coefficient is also defined as the product of temperature to the rate of change of the Seebeck coefficient with respect to temperature. So, the higher the steepness of the Seebeck coefficient in Fig. 1, the higher the value of $Th$ in magnitude as a function of $T$. Therefore, in a similar behavior to the Seebeck coefficient, the value of $Th$ is found to be more negative (i.e., more amount of heat is released) towards high temperature for EVHRG and VDWHRG. Whereas all these models overlap at lower values of temperature nearly up to $\approx 0.14$ GeV for the case of $\mu_{B}$ = 0.10 GeV. As the value of $\mu_{B}$ increases, the models start deviating from each other even at low temperatures, as in the case of the Seebeck coefficient, which is explained above.
@@
To study the efficiency of QGP medium as a thermoelectric medium, we draw the thermoelectric figure of merit, $ZT$ as a function of temperature in Fig. 4. As given in Eq. (59), $ZT$ can be defined as the ratio of $S^{2}$ to the Lorenz number ($L$), which is proportional to the ratio of thermal to electrical conductivity and defined as $L=\kappa_{0}/\sigma_{el}T$ Pradhan:2022gbm ; Singh:2023pwf . Since it is proportional to $S^{2}$, we observed that $ZT$ increases both as a function of $T$ and $\mu_{B}$. The rise in $\mu_{B}$ gives rise to $ZT$ because of the presence of more baryons as compared to antibaryons. Hence, the net thermoelectric current increases, which ultimately improves the thermoelectric efficiency of the medium. The sudden increase of ZT for the VDWHRG case can be attributed to the rise of pressure as well as number density, which subsequently affects other transport properties, due to the interplay between both attractive and repulsive interactions between the hadrons. However, the decreasing trend in the VDWHRG case may be due to the effect of liquid-gas phase transition, which comes at a lower temperature with increasing baryochemical potential.
@@ III.2 In presence of magnetic field
Here, we discuss the magneto-Seebeck coefficient, $S_{B}$ in the presence of a nonzero magnetic field as a function of $T$ for different values of $\mu_{B}$. The upper panels of Fig. 5 represent $S_{B}$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at magnetic field 0.1 $m_{\pi}^{2}$, whereas the lower panels represent those for the magnetic field 1.0 $m_{\pi}^{2}$. For the case of $eB$ = 0.1 $m_{\pi}^{2}$, we observe that $S_{B}$ increases with the temperature almost linearly up to 0.1 GeV and then starts to decrease at higher temperatures. Similar to what was observed in the case of a vanishing magnetic field, all the models agree up to a certain temperature and then start deviating from each other at higher temperatures. Furthermore, we also observe that the values of $S_{B}$ increase with the increasing values of $\mu_{B}$. In contrast to the zero magnetic field case, as shown in Eq. (87), $S_{B}$ has a dependence on the Hall-like components of electrical conductivity and other integral terms which comes into the picture because of the presence of the external magnetic field. The Hall-like component of electrical conductivity, $\sigma_{H}/T$, increases with $\mu_{B}$ because of an increase in net baryonic contributions. It is also to be noted here that there is no net contribution of mesons in the integral $\mathcal{I}_{31}/T^{2}$ and $\mathcal{I}_{42}/T^{2}$, but it enters through enthalpy per baryon ($h$). The effect of the magnetic field is significant on $\mathcal{I}_{31}/T^{2}$ at lower temperatures, but for higher temperatures, it is not effective due to the lower value of relaxation time. We also observe that $\mathcal{I}_{42}/T^{2}$ first increases with temperature and then starts to decrease. Higher is the magnetic field, and lower is the $\mathcal{I}_{42}/T^{2}$. The order of magnitude of $\sigma_{el}/T$ is larger than $\sigma_{H}/T$, whereas the order of magnitude of $\mathcal{I}_{31}/T^{2}$ is close to $\mathcal{I}_{42}/T^{2}$. Hence, the behavior of the magneto-Seebeck coefficient can be approximated as the ratio of $(\mathcal{I}_{31}/T^{2})/(\sigma_{el}/T)$ which can be seen in Fig. 5. The dependency of $\sigma_{el}/T$ on $T$ and $\mu_{B}$ in the presence of a magnetic field results in first increase of $S_{B}$ at low $T$ and then decrease at high $T$. This is because, in the presence of a magnetic field, the higher value of $\Omega_{c_{i}}\tau_{R_{i}}$ at the lower temperature range affects the mesonic contribution to $\sigma_{el}/T$ significantly. This decrease in mesonic contribution gets compensated for by an increase in baryon contribution with increasing $\mu_{B}$. On the other hand, this decrease in the mesonic contribution is not compensated by those increasing baryonic contributions for higher temperature ranges. Hence, $\sigma_{el}/T$ shows decreasing trend with increasing $\mu_{B}$ for higher temperatures. For a higher value of the magnetic field as shown in the lower panels in Fig. 5, we observe a minimum for $S_{B}$ while going to the high $\mu_{B}$ region. This minimum shifts towards high $T$ when $\mu_{B}$ increases further.
@@
Finally, we discuss the variation of the figure of merit with temperature and baryon chemical potential in the presence of an external magnetic field. Fig. 7 shows $ZT$ as a function of $T$ for three different values of $\mu_{B}$ at 0.10, 0.30, and 0.40 GeV at two different magnetic field orders, $eB$ = 0.1 $m_{\pi}^{2}$ in the upper panels, whereas the lower panels represent those for $eB$ = 1.0 $m_{\pi}^{2}$. In the presence of a magnetic field, the behavior of $ZT$ is different at low temperatures as compared to those in the case of vanishing magnetic fields. At higher temperatures, the results are nearly the same as in the case of the vanishing magnetic field, as shown in Fig. 4. As the value of $\mu_{B}$ increases, a minimum appears at low temperatures. For the case of $eB$ = 1.0 $m_{\pi}^{2}$, $ZT$ gets fully affected throughout the temperature range considered. This is because $ZT$ is proportional to the square of the magneto-Seebeck coefficient, and hence, the effect of the magnetic field on $S_{B}$ is reflected in $ZT$. The presence of magnetic field results in the highly non-monotonic behavior of $ZT$ as a function of both $T$ and $\mu_{B}$. We observe that the values of the thermoelectric figure of merit overlap for all the models at low temperatures, but a large deviation is observed at high temperatures. Along with the Ohmic-like component of thermal conductivity ($\kappa_{0}$), the Hall-like component of thermal conductivity ($\kappa_{H}$ingh:2023pwf also contributes to $ZT$ in the presence of an external magnetic field. However, the effect of $\kappa_{H}$ becomes more significant in the presence of strong magnetic field.